# Detecting High-Potential SMEs with Heterogeneous Graph Neural Networks

### v1

@@ Abstract
Small and Medium Enterprises (SMEs) constitute 99.9% of U.S. businesses and generate 44% of economic activity, yet systematically identifying high-potential SMEs remains an open challenge. We introduce SME-HGT, a Heterogeneous Graph Transformer framework that predicts which SBIR Phase I awardees will advance to Phase II funding using exclusively public data. We construct a heterogeneous graph with 32,268 company nodes, 124 research topic nodes, and 13 government agency nodes connected by approximately 99,000 edges across three semantic relation types. SME-HGT achieves an AUPRC of $0.621\pm 0.003$ on a temporally-split test set, outperforming an MLP baseline ($0.590\pm 0.002$) and R-GCN ($0.608\pm 0.013$) across five random seeds. At a screening depth of 100 companies, SME-HGT attains 89.6% precision with a $2.14\times$ lift over random selection. Our temporal evaluation protocol prevents information leakage, and our reliance on public data ensures reproducibility. These results demonstrate that relational structure among firms, research topics, and funding agencies provides meaningful signal for SME potential assessment, with implications for policymakers and early-stage investors.
@@ I Introduction
Small and Medium Enterprises (SMEs) are the primary engine of innovation and economic growth in the United States. According to the U.S. Small Business Administration [17], SMEs account for 99.9% of all U.S. businesses, employ 46.4% of the private workforce, and generate approximately 44% of U.S. economic activity. Identifying which of these firms have the highest growth potential is a longstanding challenge for policymakers, investors, and program administrators alike.
The Small Business Innovation Research (SBIR) program, established in 1982, is the U.S. government’s largest source of early-stage technology funding for small businesses, distributing over \$4 billion annually across 11 federal agencies [10]. The program operates in phases: Phase I awards (\$50K–\$275K) fund feasibility studies, while Phase II awards (\$500K–\$1.5M) support full research and development. Progression from Phase I to Phase II is competitively evaluated and serves as a meaningful proxy for a firm’s technical and commercial potential [11, 1]. Predicting which Phase I awardees will advance to Phase II is therefore both a practical policy question and a test case for data-driven firm assessment.
Current approaches to SME assessment rely predominantly on tabular features—financial metrics, patent counts, employee size—or manual expert review. These methods miss an important dimension: the relational structure in which firms are embedded. A company’s potential may be signaled not only by its own attributes but also by its connections to specific research domains, its funding relationships with government agencies, and its proximity to other successful firms in the innovation ecosystem.
Graph Neural Networks (GNNs) offer a natural framework for capturing such relational information. Recent work has demonstrated the power of GNNs for tasks ranging from social network analysis [7] to molecular property prediction [6]. In the economic domain, heterogeneous graph methods have shown promise for firm-level prediction tasks [8], though prior work has typically relied on proprietary data sources.
@@
    Public-data heterogeneous graph. We construct a heterogeneous graph of the SBIR ecosystem using exclusively public data from sbir.gov, with three node types (companies, research topics, government agencies) and three edge types capturing operational, funding, and co-topic relationships.
@@ II-A Graph Neural Networks for Economic Applications
In the context of firm-level prediction, GNNs can leverage corporate networks, ownership structures, and industry co-membership graphs to augment tabular features. However, most economic applications of GNNs operate on homogeneous graphs and rely on proprietary data. Our work addresses both limitations: we use a heterogeneous graph that explicitly models different entity types and their distinct relations, and we rely exclusively on publicly available SBIR program data.
II-B SBIR Program Analytics
The SBIR program has been extensively studied in the innovation economics literature. Lerner [10] conducted a foundational analysis showing that SBIR awardees exhibit significantly higher growth rates than matched non-awardees. Link and Scott [11] examined the program’s role in bridging the “valley of death” between basic research and commercialization. Audretsch et al. [1] studied the relationship between SBIR funding and small firm innovation output. Toole and Czarnitzki [16] analyzed Phase I to Phase II transitions specifically in the biomedical sector.
Despite this rich literature, existing SBIR analyses are predominantly retrospective and descriptive. To our knowledge, no prior work has applied graph-based machine learning to predict Phase II progression using the relational structure of the SBIR ecosystem.
@@ II-C Heterogeneous Graph Neural Networks
We adopt HGT as our primary architecture due to its flexibility in handling arbitrary heterogeneous schemas and its ability to learn type-dependent attention patterns, which we hypothesize are important for capturing the distinct roles of companies, research topics, and government agencies in the SBIR ecosystem.
@@ III-A Data
We use publicly available data from the SBIR/STTR Awards Database (data.www.sbir.gov), which contains records of all SBIR and STTR awards made by participating federal agencies. The raw dataset comprises 207,726 award records spanning multiple decades. Each record includes the company name, award amount, award year, program phase (I or II), awarding agency, and research topic code.
We apply the following data cleaning pipeline: (1) standardize company names by converting to uppercase, removing legal suffixes (Inc., LLC, Corp., etc.), and collapsing whitespace for entity resolution; (2) parse phase numbers from heterogeneous string formats; (3) extract research topic prefixes and agency identifiers; and (4) remove records with missing critical fields. After entity resolution, the cleaned dataset contains 32,268 unique companies.
@@ III-B Graph Construction
(Company, awarded_by, GovAgency)
@@
Government agency features (3 dimensions): log award count, log total funding amount, and log average award size, all normalized.
Edge construction. Each company is linked to its primary (modal) research topic via an operates_in edge and to its primary awarding agency via an awarded_by edge. Companies sharing the same primary research topic are connected by co_topic edges, with a cap of 50 edges per topic group and 20 edges per node to manage density in large topic clusters.
@@ III-C Label Design
$y_{c}=\mathbb{1}\!\left[\exists\text{ Phase~II award for }c\text{ with year}\leq t^{(1)}_{c}+5\right]$ (1)
@@ III-D SME-HGT Architecture
Input projection. Each node type $\tau\in\{\text{Company},\text{Topic},\text{GovAgency}\}$ receives a dedicated linear transformation $\mathbf{h}^{(0)}_{v}=\mathbf{W}_{\tau}\mathbf{x}_{v}+\mathbf{b}_{\tau}$, mapping heterogeneous input features to a shared hidden dimension $d=128$.
@@ IV-D Analysis
Graph structure improves discrimination. The consistent AUPRC gap between graph-based methods (HGT, R-GCN) and the non-graph MLP baseline confirms that relational structure in the SBIR ecosystem carries predictive signal beyond what tabular company features alone capture. The company-topic and company-agency relations, as well as co-topic links between firms, each contribute information about a firm’s innovation ecosystem context.
@@
Moderate overall discrimination. We note that absolute AUROC values (0.60–0.65) indicate moderate rather than strong discriminative power. This reflects the inherent difficulty of predicting Phase II outcomes—which depend on proposal quality, reviewer composition, and evolving agency priorities—from structural and historical features alone. Nevertheless, the ranking metrics demonstrate substantial practical value: screening the top 100 candidates yields nearly $90\%$ precision, more than double the base rate.
@@ V-A Policy Implications
Our results have direct relevance for SBIR program administration and small business policy. A screening tool achieving 89.6% precision among its top 100 predictions could substantially reduce the manual review burden for program managers. At a base rate of 41.8%, random selection would yield approximately 42 successful firms out of 100 reviewed; SME-HGT identifies approximately 90. This $2.14\times$ lift translates to more efficient allocation of expert review resources and potentially earlier identification of high-potential firms for targeted technical assistance.
The public-data-only design is a deliberate choice: any jurisdiction with structured grant or subsidy data can construct similar heterogeneous graphs to assess participant potential without requiring proprietary databases. This makes our approach replicable for analogous programs in the EU (Horizon Europe), UK (Innovate UK), and other national innovation systems.
@@ V-B Limitations
Several limitations warrant discussion. First, our graph relies exclusively on SBIR/STTR program data; incorporating additional public sources such as patent filings, SEC disclosures, or Census Bureau data could strengthen both node features and graph connectivity. Second, we observe only Phase I survivors—companies that already won competitive federal awards—introducing selection bias relative to the broader SME population. Third, the moderate overall discrimination ($\text{AUROC}\approx 0.65$) suggests that important predictive signals—proposal text quality, team composition, technology readiness level—are not captured in our current structural features. Fourth, our entity resolution relies on rule-based string matching, which may fail on name variants or subsidiaries not covered by our normalization rules. Finally, our evaluation uses a single temporal split; additional evaluation across different time windows would strengthen robustness claims.
@@ V-C Future Work
Several directions merit exploration: (1) incorporating textual features from award abstracts via language model embeddings to capture proposal quality signals; (2) adding temporal dynamics through time-aware graph architectures such as evolving GCNs; (3) conducting ablation studies to quantify the marginal contribution of each node and edge type; (4) extending to downstream prediction targets such as patent filing, revenue growth, or acquisition events; and (5) developing an interactive dashboard for SBIR program managers to explore model predictions alongside company profiles.
@@ VI Conclusion
We presented SME-HGT, a Heterogeneous Graph Transformer framework for identifying high-potential small businesses using public SBIR program data. By constructing a heterogeneous graph capturing relationships among 32,268 companies, 124 research topics, and 13 government agencies, we demonstrated that relational structure provides meaningful predictive signal beyond tabular features alone. SME-HGT achieves an AUPRC of 0.621 on temporally-split test data, outperforming both MLP (+3.1+3.1 pp) and R-GCN (+1.3+1.3 pp) baselines while maintaining low variance across seeds. At a practical screening depth of 100 candidates, the model attains 89.6% precision—a $2.14\times$ improvement over random selection. Our strict temporal evaluation protocol and exclusive reliance on public data ensure that these results reflect realistic deployment conditions. This work illustrates the potential of heterogeneous graph neural networks for policy-relevant prediction tasks in the SME innovation ecosystem.
@@ Acknowledgments
Data sourced from the SBIR/STTR Awards Database maintained by the U.S. Small Business Administration. Computational experiments were conducted using PyTorch and PyTorch Geometric.

### v2

# arXiv:2602.19591v2[cs.LG] 27 Feb 2026

@@ Abstract
Small and Medium Enterprises (SMEs) constitute the vast majority of businesses globally and generate a significant portion of economic activity, yet systematically identifying high-potential SMEs remains an open challenge. We introduce SME-HGT, a Heterogeneous Graph Transformer framework that predicts which early-stage (Phase I) innovation grant awardees will advance to Phase II funding using exclusively public data. We construct a heterogeneous graph with 32,268 company nodes, 124 research topic nodes, and 13 funding agency nodes connected by approximately 99,000 edges across three semantic relation types. SME-HGT achieves an AUPRC of $0.621\pm 0.003$ on a temporally-split test set, outperforming an MLP baseline ($0.590\pm 0.002$) and R-GCN ($0.608\pm 0.013$) across five random seeds. At a screening depth of 100 companies, SME-HGT attains 89.6% precision with a $2.14\times$ lift over random selection. Our temporal evaluation protocol prevents information leakage, and our reliance on public data ensures reproducibility. These results demonstrate that relational structure among firms, research topics, and funding bodies provides meaningful signal for SME potential assessment, with implications for economic policymakers and early-stage investors globally.
@@ I Introduction
Small and Medium Enterprises (SMEs) are the primary engine of innovation and economic growth globally. According to international economic development reports [18], SMEs account for the overwhelming majority of all businesses, employ a large segment of the private workforce, and generate a substantial portion of global economic activity. Identifying which of these firms have the highest growth potential is a longstanding challenge for economic analysts, investors, and grant program administrators alike.
National and regional innovation grant programs are major sources of early-stage technology funding for small businesses, distributing billions annually across various funding agencies [11]. These competitive programs typically operate in sequential phases: Phase I awards fund initial feasibility studies, while Phase II awards support full research and development. Progression from Phase I to Phase II is rigorously evaluated and serves as a meaningful proxy for a firm’s technical and commercial potential [12, 1]. Predicting which Phase I awardees will advance to Phase II is therefore both a practical policy question and a test case for data-driven firm assessment.
Current approaches to SME assessment rely predominantly on tabular features—financial metrics, patent counts, employee size—or manual expert review. These methods miss an important dimension: the relational structure in which firms are embedded. A company’s potential may be signaled not only by its own attributes but also by its connections to specific research domains, its funding relationships with sponsoring agencies, and its proximity to other successful firms in the innovation ecosystem.
Graph Neural Networks (GNNs) offer a natural framework for capturing such relational information. Recent work has demonstrated the power of GNNs for tasks ranging from social network analysis [8] to molecular property prediction [7]. Broader artificial intelligence methodologies are similarly transforming critical physical infrastructure, optimizing design and control strategies in power electronics and energy systems [4]. In the economic domain, heterogeneous graph methods have shown promise for firm-level prediction tasks [9], though prior work has typically relied on proprietary data sources.
@@
    Public-data heterogeneous graph. We construct a heterogeneous graph of an innovation grant ecosystem using exclusively public database records, with three node types (companies, research topics, funding agencies) and three edge types capturing operational, funding, and co-topic relationships.
@@ II-A Graph Neural Networks for Economic Applications
In the context of firm-level prediction, GNNs can leverage corporate networks, ownership structures, and industry co-membership graphs to augment tabular features. However, most economic applications of GNNs operate on homogeneous graphs and rely on proprietary data. Our work addresses both limitations: we use a heterogeneous graph that explicitly models different entity types and their distinct relations, and we rely exclusively on publicly available grant program data.
II-B Innovation Grant Analytics
Innovation grant programs have been extensively studied in the economics literature. Lerner [11] conducted a foundational analysis showing that grant awardees exhibit significantly higher growth rates than matched non-awardees. Link and Scott [12] examined the role of these programs in bridging the “valley of death” between basic research and commercialization. Audretsch et al. [1] studied the relationship between early-stage funding and small firm innovation output. Toole and Czarnitzki [17] analyzed phase progression specifically in the biomedical sector.
Despite this rich literature, existing grant analyses are predominantly retrospective and descriptive. To our knowledge, no prior work has applied graph-based machine learning to predict Phase II progression using the relational structure of the innovation ecosystem.
@@ II-C Heterogeneous Graph Neural Networks
We adopt HGT as our primary architecture due to its flexibility in handling arbitrary heterogeneous schemas and its ability to learn type-dependent attention patterns, which we hypothesize are important for capturing the distinct roles of companies, research topics, and funding agencies.
@@ III-A Data
We use publicly available data from an open innovation awards database, which contains records of public grants made by participating funding bodies. The raw dataset comprises 207,726 award records spanning multiple decades. Each record includes the company name, award amount, award year, program phase (I or II), awarding agency, and research topic code.
We apply the following data cleaning pipeline: (1) standardize company names by converting to uppercase, removing legal suffixes, and collapsing whitespace for entity resolution; (2) parse phase numbers from heterogeneous string formats; (3) extract research topic prefixes and agency identifiers; and (4) remove records with missing critical fields. After entity resolution, the cleaned dataset contains 32,268 unique companies.
@@ III-B Graph Construction
(Company, awarded_by, FundingAgency)
@@
Funding agency features (3 dimensions): log award count, log total funding amount, and log average award size, all normalized.
Edge construction. Each company is linked to its primary (modal) research topic via an operates_in edge and to its primary funding body via an awarded_by edge. Companies sharing the same primary research topic are connected by co_topic edges, with a cap of 50 edges per topic group and 20 edges per node to manage density in large topic clusters.
@@ III-C Label Design
$y_{c}=\mathbb{1}\!\left[\exists\text{ Phase II award for }c\text{ with year}\leq t^{(1)}_{c}+5\right]$ (1)
@@ III-D SME-HGT Architecture
Input projection. Each node type $\tau\in\{\text{Company},\text{Topic},\text{FundingAgency}\}$ receives a dedicated linear transformation $\mathbf{h}^{(0)}_{v}=\mathbf{W}_{\tau}\mathbf{x}_{v}+\mathbf{b}_{\tau}$, mapping heterogeneous input features to a shared hidden dimension $d=128$.
@@ IV-D Analysis
Graph structure improves discrimination. The consistent AUPRC gap between graph-based methods (HGT, R-GCN) and the non-graph MLP baseline confirms that relational structure in the innovation ecosystem carries predictive signal beyond what tabular company features alone capture. The company-topic and company-agency relations, as well as co-topic links between firms, each contribute information about a firm’s context.
@@
Moderate overall discrimination. We note that absolute AUROC values (0.60–0.65) indicate moderate rather than strong discriminative power. This reflects the inherent difficulty of predicting Phase II outcomes—which depend on proposal quality, reviewer composition, and evolving funding priorities—from structural and historical features alone. Nevertheless, the ranking metrics demonstrate substantial practical value: screening the top 100 candidates yields nearly $90\%$ precision, more than double the base rate.
@@ V-A Policy Implications
Our results have direct relevance for global grant program administration and small business policy. A screening tool achieving 89.6% precision among its top 100 predictions could substantially reduce the manual review burden for program managers. At a base rate of 41.8%, random selection would yield approximately 42 successful firms out of 100 reviewed; SME-HGT identifies approximately 90. This $2.14\times$ lift translates to more efficient allocation of expert review resources and potentially earlier identification of high-potential firms for targeted technical assistance.
The public-data-only design is a deliberate choice: any jurisdiction with structured grant or subsidy data can construct similar heterogeneous graphs to assess participant potential without requiring proprietary databases. This makes our approach replicable for analogous programs across various national innovation systems.
@@ V-B Limitations
Several limitations warrant discussion. First, our graph relies exclusively on historical grant program data; incorporating additional public sources such as patent filings or local economic data could strengthen both node features and graph connectivity. Second, we observe only Phase I survivors—companies that already won competitive grants—introducing selection bias relative to the broader SME population. Third, the moderate overall discrimination ($\text{AUROC}\approx 0.65$) suggests that important predictive signals—proposal text quality, team composition, technology readiness level—are not captured in our current structural features. Fourth, our entity resolution relies on rule-based string matching, which may fail on name variants or subsidiaries not covered by our normalization rules. Finally, our evaluation uses a single temporal split; additional evaluation across different time windows would strengthen robustness claims.
@@ V-C Future Work
Several directions merit exploration: (1) incorporating textual features from award abstracts via language model embeddings to capture proposal quality signals; (2) adding temporal dynamics through time-aware graph architectures such as evolving GCNs; (3) conducting ablation studies to quantify the marginal contribution of each node and edge type; (4) extending to downstream prediction targets such as patent filing, revenue growth, or acquisition events; and (5) developing an interactive dashboard for program managers to explore model predictions alongside company profiles.
@@ VI Conclusion
We presented SME-HGT, a Heterogeneous Graph Transformer framework for identifying high-potential small businesses using public grant program data. By constructing a heterogeneous graph capturing relationships among 32,268 companies, 124 research topics, and 13 funding agencies, we demonstrated that relational structure provides meaningful predictive signal beyond tabular features alone. SME-HGT achieves an AUPRC of 0.621 on temporally-split test data, outperforming both MLP (+3.1+3.1 pp) and R-GCN (+1.3+1.3 pp) baselines while maintaining low variance across seeds. At a practical screening depth of 100 candidates, the model attains 89.6% precision—a $2.14\times$ improvement over random selection. Our strict temporal evaluation protocol and exclusive reliance on public data ensure that these results reflect realistic deployment conditions. This work illustrates the potential of heterogeneous graph neural networks for policy-relevant prediction tasks in the global SME innovation ecosystem.
@@ Acknowledgments
Data sourced from public open-data grant portals. Computational experiments were conducted using PyTorch and PyTorch Geometric.